# The contact process on dynamical random trees with degree dependence

### v1

@@ 3.1 Heuristics for the results on general evolving graphs
The formal proof proceeds as follows. After coupling these two processes we use a submartingale argument for a specifically chosen function of the wait-and-see process, which takes the number of revealed edges into account. It is shown that under the assumptions stated in Theorem 3.1, there exists a sufficiently small $\lambda$ such that the wait-and-see process dies out almost surely.
@@ 3.2 Heuristic for the evolving tree
Thus, once an order of $N^{1-\alpha-\eta}$ many infected neighbours is reached, the total number of infected neighbours will be kept at this order for a time of order $S=e^{cN^{1-\alpha-2\eta}}$ where $c>0$. This implies that if $\alpha+2\eta<1$, then the infection persists for an exponentially long time with high probability. This will then be enough time to pass the infection on to the next star.
@@
for some $10<K<$ 1 (see Lemma 7.7 for details). By the previous heuristics we know that an infection can survive in a star of size $N$ for a time of order $S=e^{cN^{1-\alpha-2\eta}}$, which means that for a time period of length $S$ it is possible for the infection to be transmitted from $\rho$ to $z$ without the infection in the star around $\rho$ becoming extinct. Thus, the probability of a successful transmission can be lower bounded by roughly $S/(4r)$ trials of attempting to infect vertex $z$ in one go in a time span of length less than $4r$. Then the probability that at least one trial is successful is approximately
@@
$-\log\mathbb{P}(D_{\rho}=N)\ll r_{N}\ll c\lambda^{2}N^{1-\alpha-2\eta}$ (3.12)
@@
On the other hand, if $\alpha\in\big{(}\tfrac{1-2\eta}{2},1\big{)}$ then this argument is no longer true for all $\lambda>0$. But it can be modified such that we can conclude strong survival if the infection rate $\lambda$ is chosen large enough. First, the size $N$ of the stars is chosen large but fixed, then by (3.12) if $\lambda$ is chosen to be at least of order $N^{2\eta}$ it still holds true that the infection is transmitted back and forth between stars of size $N$ with a sufficiently high probability. This can then be used to show $(ii)$, i.e. strong survival for a sufficiently large $\lambda$ as long as $\alpha<1$.
@@
Our proof strategy is similar to the strategy proposed for the static case by Pemantle [18]. But there are some significant differences. The basic idea for the survival of the infection in a star is that if there is a sufficient number of infected neighbours, in our case $\lambda N^{1-\alpha-\eta}$, then the infection can persist on its own for a long time. In comparison to the static case we lose roughly $N^{-\eta}$ many infections in the first steps since additionally to recoveries edges update. Since the star is exceptionally large the edge will be most likely closed after an update. Thus, only $\lambda N^{1-\alpha-\eta}$ many neighbours can be sustained and we lose another $N^{\eta}$ for the same reason in the probability of reinfection, which is $1-\exp(\lambda^{2}N^{1-\alpha-2\eta})$. In the static case, we would have in both terms $N^{1-\alpha}$ instead.
@@ 4 Discussion
In this section we discuss our main results in the context of the CPDG on a BGW tree. We concentrate on our main example, i.e. we assume that the update speed $v$ satisfies (2.3) and the connection probability is of the form $p=p_{\alpha,\sigma}$ as defined in (2.5). This means that for every $m\leq n$ there exists an $\nu_{1},\nu_{2}>0$ such that
@@
where $\kappa,\alpha>0$, $\eta\in\mathbb{R}$ and $\sigma\in[0,1]$.
@@ 4.1 Discussion for power law offspring distributions
Consider an offspring distribution $\zeta$ that follows a power law, i.e. $\mathbb{P}(\zeta=N)\sim cN^{-{b}}$ for $N$ large with ${b}>1$. An immediate consequence of Corollary 3.2, Theorem 3.4 and Proposition 3.5 is the following result.
@@
    If either $\eta\leq 0$ and $\alpha\in[0,1)$ or if $\eta\leq 1/2$ and $\alpha\in(0,1-2\eta)$ then $\lambda_{1}=\lambda_{2}=0$.
@@
    If $\eta\geq 0$, $01-2\alpha>$ 0 and $\mathbb{P}(\zeta=0)=0$ then $\lambda_{1}=\lambda_{2}=0$.
@@
We illustrate the resulting phase diagram in Figure 2. A first observation is that the choice of $\kappa$,$\nu_{1},\nu_{2}$ as well as the offspring tail exponent $b$ have no effect on the existence or non-existence of a subritical phase.
@@
It is plausible to assume that if we do not allow any updates in the tree, then the contact process on this tree behaves in the same way as the contact process on a static BGW($\zeta^{p}$) tree. The difference coming from the correlation of the percolation probabilities of consecutive generations should be negligible. Assuming that $\alpha<1$ and that $\zeta$ has a power law distribution with exponent $b$. Then, using a local CLT for binomial random variables and their strong concentration properties, one can show that
$\mathbb{P}(\zeta^{p}=k)&\asymp\sum_{n=k}^{\infty}n^{-b}\mathbb{P}(\textrm{Bin}(n,n^{-\alpha})=k)\\ &\asymp\sum_{n=k}^{\infty}n^{-b}\frac{1}{\sqrt{n^{1-\alpha}}}e^{-\frac{(k-n^{1-\alpha})^{2}}{2n^{1-\alpha}(1-n^{-\alpha})}}\\ &\asymp k^{-\frac{b-\alpha}{1-\alpha}}.\end{aligned}.$
where $\asymp$ is to be interpreted as upper and lower bounds that agree up to multiplicative constants and we use that the sum concentrates around those values of $n$ such that $n^{1-\alpha}-\sqrt{n^{1-\alpha}}\leq k\leq n^{1-\alpha}+\sqrt{n^{1-\alpha}}$. In particular, $\zeta^{p}$ has a power law distribution with exponent $(b-\alpha)/(1-\alpha)$. The results for the contact process on a static BGW-tree therefore suggest that in the zero update speed case, we always have that for the corresponding critical value $\lambda^{\rm static}_{2}=0$. In the case $\alpha\geq 1$, a standard calculation shows that the percolated offspring distribution $\zeta^{p}$ has bounded exponential moments, so on a static BGW$(\zeta^{p})$-tree the contact process has a phase transition.
Indeed, comparing to our model, we have that in the slow speed regime when $\eta\leq 0$ and $\alpha<1$, then our model exhibits a phase transition in the same cases as the model on a static BGW($\zeta^{p}$)-tree.
@@
Explosion. Finally, we want to comment on the possibility that the CPDG started with finitely many infected vertices explodes in finite time with positive probability. If the offspring distribution $\zeta$ has finite mean, i.e. $\mathbb{E}[\zeta]<\infty$, then the process does not explode in finite time almost surely. This follows by a direct coupling with the classical contact process, where it is know that the process does almost surely not explode if $\mathbb{E}[\zeta]<\infty$, see e.g.  [4, Theorem 3.2] in the case when the fitness is constant equal to one.
@@ 4.2 Discussion of stretched exponential offspring distributions
    If $\eta\geq 0$, $01-\beta-2\alpha>$ 0 and $\mathbb{P}(\zeta=0)=0$, then $\lambda_{1}=\lambda_{2}=0$.
@@
In the case $\alpha\geq 1$ our results do not change, which is not surprising since these results do not depend on the specific graph structure, but only how $p$ depends on the degrees. But if $\alpha<1$, choosing this class of distributions with slightly “lighter” tails compared to the power law distributions has in fact consequences for Theorem 3.4 and Proposition 3.5. In this case we get an additional restriction for our parameter regime, in that $\alpha+2\eta<1-\beta$ needs to be satisfied for (3.7) to hold so that we can deduce $\lambda_{2}=0$ from Theorem 3.4$(i)$. Similary, the condition $\alpha<1-\beta$ is necessary so that (3.8) holds and so that Theorem 3.4$(ii)$ yields $\lambda_{2}<\infty$. We illustrate the additional restrictions in Figure 3 with the product kernel $p_{\alpha,1}$ as connection probability.
@@
For the fast update speed regime, i.e. $\eta$ sufficiently large, the comparison with the penalised contact process yields that for $\alpha<(1-\beta)/2$ we have $\lambda_{2}=0$. However, for the penalised contact process, as well as for our model, there is a gap in the fast speed case as illustrated by the grey area in Figure 3, where it is not clear whether a phase transition exists or not. Similarly to the case of power-law tails, we would expect a linear transition from the static to the penalised behaviour. However, if we consider the heuristic explanation in Subsection 3.2 and see in the last part that if $\alpha+2\eta>1-\beta$, then the probability of finding another star of the same size in a reasonable number of steps is not necessarily bounded away from 00 such that the strategy of survival through stars might not work.
@@
$\limsup_{N\to\infty}\frac{\log\mathbb{P}(\zeta=N)}{N^{1-\alpha-2(\eta\vee 0)}}\in(-\infty,0)$
@@
Assume $p$ and $v$ satisfy (4.1) and that $\alpha\geq 1-\beta$. For $\kappa>0$ small enough, does there exist a $\nu^{\prime}=\nu^{\prime}(\kappa,\eta)$ such that $\lambda_{1}=\infty$ for all $\nu_{2}<\nu^{\prime}$
@@ 5.1 Graphical representation
for all $t\in[T_{n}(e),T_{n+1}(e))$ with $n\geq 1$. Then we define $\mathbf{B}_{t}:=\{e\in\mathcal{E}:X_{t}(e)=1\}$.
Finally, we define the infection process $\mathbf{C}$. Let $\Delta^{\inf}=(\Delta^{\inf}_{\{x,y\}})_{\{x,y\}\in E}$ and $\Delta^{\text{rec}}=(\Delta^{\text{rec}}_{x})_{x\in V}$ be two independent families of Poisson processes such that for all $\{x,y\}\in E$ and $x\in V$ fixed, the processes $\Delta^{\inf}_{\{x,y\}}$ and $\Delta^{\text{rec}}_{x}$ have rate $\lambda$ and $1$, respectively. Now we need to introduce the notion of an infection path. Let $(y,s)$ and $(x,t)$ with $s<t$ be space-time points. We write $(y,s)\stackrel{{\scriptstyle\mathbf{B}}}{{\longrightarrow}}(x,t)$ if there exists a sequence of times $s=t_{0}<t_{1}<\dots<t_{n}\leq t_{n+1}=t$ and space points $y=x_{0},x_{1},\dots,x_{n}=x$ such that ($(\{x_{k},x_{k+1}\},t_{k})\in\Delta^{\inf}$ as well as $\{x_{k},x_{k+1}\}\in\mathbf{B}_{t_{k}}$ and $\Delta^{\text{rec}}\cap\big{(}\{x_{k}\}\times[t_{k},t_{k+1})\big{)}=\emptyset$ for all $k\in\{0,\dots,n\}$ with $n\in\mathbb{N}$.
@@ 5.2 Properties of $(\mathbf{C},\mathbf{B})$ on $\mathcal{T}$
$\lambda_{1}(\mathcal{T})=\lambda_{1}=\inf\big{\{}\lambda>0:\mathbb{P}^{\{\rho\}}(\mathbf{C}_{t}\neq\emptyset\,\forall\,t\geq 0)>0\big{\}}\\ \lambda_{2}(\mathcal{T})=\lambda_{2}=\inf\big{\{}\lambda>0:\liminf_{t\to\infty}\mathbb{P}^{\{\rho\}}(\rho\in\mathbf{C}_{t})>0\big{\}}\end{aligned}$
on the event $\{|\mathcal{T}|=\infty\}$.
The proof follows similar ideas as those used first in [18, Proposition 3.1] and afterwards in [8, Proposition 1].
@@
$\mathbb{P}\big{(}\theta_{\mathcal{T}}(\lambda,v,\{\rho\})=0\mid|\mathcal{T}|=\infty\big{)}\in\{0,1\}$
because if $|\mathcal{T}|<\infty$ then $\theta_{\mathcal{T}}(\lambda,v,\{\rho\})=0$, and so $\mathbb{P}\big{(}\theta_{\mathcal{T}}(\lambda,v,\{\rho\})=0,|\mathcal{T}|<\infty\big{)}=\mathbb{P}\big{(}|\mathcal{T}|<\infty\big{)}$. Now, assume that the root $\rho$ has $k$ children, namely $x_{1},\dots,x_{k}$. Then the subtrees rooted in these children are independent identically distributed BGW trees with again offspring distribution $\zeta$, which we denote by $\mathcal{T}_{1},\dots,\mathcal{T}_{k}$. From Lemma 7.6 below, we know that the probability to push the infection from $\rho$ to some $x_{i}$ conditioned on $\{\rho,x_{i}\}\notin\mathbf{B}_{0}$ is given by
$b_{\rho,x_{i}}:=\frac{\lambda v(D_{\rho},D_{x_{i}})p(D_{\rho},D_{x_{i}})}{\lambda+v(D_{\rho},D_{x_{i}})+\lambda v(D_{\rho},D_{x_{i}})p(D_{\rho},D_{x_{i}})+1}.$
Note that by monotonicity
@@
It follows that if  $\theta_{\mathcal{T}_{i}}(\lambda,v,\{x_{i}\})>0$ for some $i\in\{1,\dots,k\}$, then  $\theta_{\mathcal{T}}(\lambda,v,\{\rho\})>0$. In other words, if the contact process dies out on $\mathcal{T}$ then it has to die out on all subtrees $\mathcal{T}_{1},\dots,\mathcal{T}_{k}$. Denote by $g$ the generating function of the random variable $\zeta$. Now we observe that

### v2

@@ 3.1 Heuristics for the results on general evolving graphs
The formal proof proceeds as follows. After coupling the CPDG and the wait-and-see process we use a submartingale argument for a specifically chosen function of the wait-and-see process, which takes the number of revealed edges into account. It is shown that under the assumptions stated in Theorem 3.1, there exists a sufficiently small $\lambda$ such that the wait-and-see process dies out almost surely.
@@ 3.2 Heuristic for the evolving tree
Thus, once an order of $N^{1-\alpha-\eta}$ many infected neighbours is reached, the total number of infected neighbours will be kept at this order for a time of order $S=e^{c\lambda^{2}N^{1-\alpha-2\eta}}$ where $c>0$. This implies that if $\alpha+2\eta<1$, then the infection persists for a stretched exponentially long time with high probability. This will then be enough time to pass the infection on to the next star. Note that we ignore all neighbours whose shared edge with the centre of the star were updated. These are indeed negligible, in Remark 3.6 this is explained in more detail.
@@
for some $0<K<1$ (see Lemma 7.6 for details). By the previous heuristics we know that an infection can survive in a star of size $N$ for a time of order $S=e^{c\lambda^{2}N^{1-\alpha-2\eta}}$, which means that for a time period of length $S$ it is possible for the infection to be transmitted from $\rho$ to $z$ without the infection in the star around $\rho$ becoming extinct. Thus, the probability of a successful transmission can be lower bounded by roughly $S/(4r)$ trials of attempting to infect vertex $z$ in one go in a time span of length less than $4r$. Then the probability that at least one trial is successful is approximately
@@
$\log\mathbb{P}(D_{\rho}=N)\ll r_{N}\ll c\lambda^{2}N^{1-\alpha-2\eta}.$ (3.12)
@@
On the other hand, if $\alpha\in\big((1-2\eta)\vee 0,1\big)$ then this argument is no longer true for all $\lambda>0$. But it can be modified such that we can conclude strong survival if the infection rate $\lambda$ is chosen large enough. First, the size $N$ of the stars is chosen large but fixed, then by (3.12) if $\lambda$ is chosen to be at least of order $N^{2\eta}$ it still holds true that the infection is transmitted back and forth between stars of size $N$ with a sufficiently high probability. This can then be used to show $(ii)$, i.e. strong survival for a sufficiently large $\lambda$ as long as $\alpha<1$.
@@
Our proof strategy is similar to the strategy proposed for the static case by Pemantle [18]. But there are some significant differences. The basic idea for the survival of the infection in a star is that if there is a sufficient number of infected neighbours, in our case $\lambda N^{1-\alpha-\eta}$, then the infection can persist on its own for a long time. In comparison to the static case we lose roughly a factor of $N^{-\eta}$ many infections in the first step since additionally to recoveries edges update. Since the star is exceptionally large the edge will be most likely closed after an update. Thus, only $\lambda N^{1-\alpha-\eta}$ many neighbours can be sustained and we lose another factor of $N^{-\eta}$ for the same reason in the probability of reinfection, which is $1-\exp(-\lambda^{2}N^{1-\alpha-2\eta})$. In the static case, we would have in both terms $N^{1-\alpha}$ instead.
Clearly as $\eta$ increases this strategy becomes worse and is no longer applicable for $\eta\geq\tfrac{1}{2}$. On the other hand, neighbours $y$ whose shared edge has been updated after receiving the infection from $x$, infect $x$ at rate $\lambda p(D_{x},D_{y})$. This is the same rate as for the penalised contact process. So for $\eta\geq\tfrac{1}{2}$ switching to a strategy in which only updated neighbours are considered, using heuristics similar to the above, the stars of size $N$ should retain the infection for a time of order $e^{c\lambda^{2}N^{1-2\alpha}}$ with probability $1-\exp(-\lambda^{2}N^{1-2\alpha})$, which is comparable to the results shown in [1].
This suggests that in the parameter regime of moderate update speed there is a change of behaviour, such that the process behaves less like a contact process on a static pruned tree, but rather like a penalised version on the full tree. The tipping point seems to be at $\eta=\tfrac{1}{4}$.
@@ 4 Discussion
In this section we discuss our main results in the context of the CPDG on a BGW tree. We concentrate on our main example, i.e. we assume that the update speed $v$ satisfies (2.3) and the connection probability is of the form $p=p_{\alpha,\sigma}$ as defined in (2.6). This means that for every $m\in\mathbb{N}$ there exist $\nu_{1},\nu_{2}>0$ such that
@@
for all $m\leq n$, where $\kappa,\alpha>0$, $\eta\in\mathbb{R}$ and $\sigma\in[0,1]$.
@@ 4.1 Discussion for power law offspring distributions
Consider an offspring distribution $\zeta$ that follows a power law, i.e. $\mathbb{P}(\zeta=N)\sim cN^{-{b}}$ for $N$ large with b>1{b}>1. An immediate consequence of Corollary 3.2, Theorem 3.4 and Proposition 3.5 is the following result.
@@
    If either $\eta\leq 0$ and $\alpha\in[0,1)$ or if $0\leq\eta\leq 1/2$ and $\alpha\in(0,1-2\eta)$ then $\lambda_{1}=\lambda_{2}=0$.
@@
    If $\eta\geq 0$, $1-2\alpha>0$ and $\mathbb{P}(\zeta=0)=0$ then $\lambda_{1}=\lambda_{2}=0$.
@@
We illustrate the resulting phase diagram in Figure 2. A first observation is that the choice of $\kappa$$\nu_{1},\nu_{2}$ as well as the offspring tail exponent $b$ have no effect on the existence or non-existence of a subcritical phase.
@@
It is plausible to assume that if we do not allow any updates in the tree, then the contact process on this tree behaves in the same way as the contact process on a static BGW($\zeta^{p}$) tree. The difference coming from the correlation of the percolation probabilities of consecutive generations should be negligible. Assume that $\alpha<1$ and that $\zeta$ has a power law distribution with exponent $b$. Then, using a local CLT for binomial random variables and their strong concentration properties, one can show that
$\mathbb{P}(\zeta^{p}=k)$ $\asymp\sum_{n=k}^{\infty}n^{-b}\mathbb{P}(\textrm{Bin}(n,n^{-\alpha})=k)$
$\asymp\sum_{n=k}^{\infty}n^{-b}\frac{1}{\sqrt{n^{1-\alpha}}}e^{-\frac{(k-n^{1-\alpha})^{2}}{2n^{1-\alpha}(1-n^{-\alpha})}}$
$\asymp k^{-\frac{b}{1-\alpha}-\frac{1}{2}}\sum_{n:|n^{1-\alpha}-k|\leq\sqrt{k}}e^{-\frac{(k-n^{1-\alpha})^{2}}{2n^{1-\alpha}(1-n^{-\alpha})}}$
$\asymp k^{-\frac{b}{1-\alpha}-\frac{1}{2}}\Big(\big(k+\sqrt{k}\big)^{\frac{1}{1-\alpha}}-\big(k-\sqrt{k}\big)^{\frac{1}{1-\alpha}}\Big)$
$\asymp k^{-\frac{b-\alpha}{1-\alpha}}.$
[…]
Indeed, comparing to our model, we have thatwhen $\eta=0$, then our model exhibits a phase transition in the same cases as the model on a static BGW($\zeta^{p}$)-tree, i.e. when $\alpha\geq 1$. Moreover we conjecture that the same is true in the full slow speed regime $\eta\leq 0$.
@@
Explosion. Finally, we want to comment on the possibility that the CPDG started with finitely many infected vertices explodes in finite time with positive probability. If the offspring distribution $\zeta$ has finite mean, i.e. $\mathbb{E}[\zeta]<\infty$, then the process does not explode in finite time almost surely. This follows by a direct coupling with the classical contact process, where it is know that the process surely does not explode if $\mathbb{E}[\zeta]<\infty$, see e.g. [5, Theorem 3.2] in the case when the fitness is constant equal to one.
@@ 4.2 Discussion of stretched exponential offspring distributions
    If $\eta\geq 0$, $1-\beta-2\alpha>0$ and $\mathbb{P}(\zeta=0)=0$, then $\lambda_{1}=\lambda_{2}=0$.
@@
In the case $\alpha\geq 1$ our results do not change, which is not surprising since these results do not depend on the specific graph structure, but only how $p$ depends on the degrees. But if $\alpha<1$, choosing this class of distributions with slightly “lighter” tails compared to the power law distributions has in fact consequences for Theorem 3.4 and Proposition 3.5. In this case we get an additional restriction for our parameter regime, in that $\alpha+2\eta<1-\beta$ needs to be satisfied for (3.7) to hold so that we can deduce $\lambda_{2}=0$ from Theorem 3.4(i)(i). Similary, the condition $\alpha<1-\beta$ is necessary so that (3.8) holds and so that Theorem 3.4(ii)(ii) yields $\lambda_{2}<\infty$. We illustrate the additional restrictions in Figure 3 with the product kernel $p_{\alpha,1}$ as connection probability.
@@
For $\eta>0$, the comparison with the penalised contact process yields that for $\alpha<(1-\beta)/2$ we have $\lambda_{2}=0$. However, for the penalised contact process, as well as for our model, there is a gap in the fast speed case as illustrated by the grey area in Figure 3, where it is not clear whether a phase transition exists or not. Similarly to the case of power-law tails, we would expect a linear transition from the static to the penalised behaviour. However, if we consider the heuristic explanation in Subsection 3.2 and see in the last part that if $\alpha+2\eta>1-\beta$, then the probability of finding another star of the same size in a reasonable number of steps is not necessarily bounded away from 0 such that the strategy of survival through stars might not work.
@@
$\limsup_{N\to\infty}\frac{\log\mathbb{P}(\zeta=N)}{N^{1-\alpha-2(\eta\vee 0)}}\in(-\infty,0)\$
@@
Assume $p$ and $v$ satisfy (4.1) and that $\alpha\geq 1-\beta$. For $\kappa>0$ small enough, does there exist a $\nu^{\prime}=\nu^{\prime}(\kappa,\eta)$ such that $\lambda_{1}=\infty$ for all $\nu_{2}<\nu^{\prime}$?
@@ 5.1 Graphical representation
for all $t\in[T_{n}(e),T_{n+1}(e))$ with $n\geq 1$. Then we define $\mathbf{B}_{t}:=\{e\in E:X_{t}(e)=1\}$.
Finally, we define the infection process $\mathbf{C}$. Let $\Delta^{\inf}=(\Delta^{\inf}_{\{x,y\}})_{\{x,y\}\in E}$ and $\Delta^{\text{rec}}=(\Delta^{\text{rec}}_{x})_{x\in V}$ be two independent families of Poisson processes such that for all $\{x,y\}\in E$ and $x\in V$ fixed, the processes $\Delta^{\inf}_{\{x,y\}}$ and $\Delta^{\text{rec}}_{x}$ have rate $\lambda$ and $1$, respectively. Now we need to introduce the notion of an infection path. Let $(y,s)$ and $(x,t)$ with $s<t$ be space-time points. We write $(y,s)\stackrel{{\scriptstyle\mathbf{B}}}{{\longrightarrow}}(x,t)$ if there exists a sequence of times $s=t_{0}<t_{1}<\dots<t_{n}\leq t_{n+1}=t$ and space points $y=x_{0},x_{1},\dots,x_{n}=x$ such that $(\{x_{k},x_{k+1}\},t_{k})\in\Delta^{\inf}$ as well as $\{x_{k},x_{k+1}\}\in\mathbf{B}_{t_{k}}$ and $\Delta^{\text{rec}}\cap\big(\{x_{k}\}\times[t_{k},t_{k+1})\big)=\emptyset$ for all $k\in\{0,\dots,n\}$ with $n\in\mathbb{N}$.
@@ 5.2 Properties of $(\mathbf{C},\mathbf{B})$ on $\mathcal{T}$
$\begin{aligned} \lambda_{1}(\mathcal{T})&=\lambda_{1}:=\inf\big\{\lambda>0:\mathbb{P}^{\{\rho\}}(\mathbf{C}_{t}\neq\emptyset\,\forall\,t\geq 0)>0\big\}\\ \lambda_{2}(\mathcal{T})&=\lambda_{2}:=\inf\big\{\lambda>0:\mathbb{P}^{\{\rho\}}(\forall s\geq 0\ \exists t\geq s\,:\,\rho\in\mathbf{C}_{t})>0\big\}\end{aligned}$
almost surely on the event $\{|\mathcal{T}|=\infty\}$.
The proof follows similar ideas as those used first in [18, Proposition 3.1] and afterwards in [9, Proposition 1].
@@
$\mathbb{P}\big(\theta_{\mathcal{T}}(\lambda,v,\{\rho\})=0\mid|\mathcal{T}|=\infty\big)\in\{0,1\}$ (5.3)
because if $|\mathcal{T}|<\infty$ then $\theta_{\mathcal{T}}(\lambda,v,\{\rho\})=0$, and so $\mathbb{P}\big(\theta_{\mathcal{T}}(\lambda,v,\{\rho\})=0,|\mathcal{T}|<\infty\big)=\mathbb{P}\big(|\mathcal{T}|<\infty\big)$. Now, assume that the root $\rho$ has $k$ children, namely $x_{1},\dots,x_{k}$. Then the subtrees rooted in these children are independent identically distributed BGW trees with again offspring distribution $\zeta$, which we denote by $\mathcal{T}_{1},\dots,\mathcal{T}_{k}$. Let $b_{\rho,x_{i}}$ be the probability that $\rho$ infects $x_{i}$ conditional on $\{\rho,x_{i}\}\notin\mathbf{B}_{0}$. Note that by monotonicity
@@
As $b_{\rho,x_{i}}>0$, it follows that if $\theta_{\mathcal{T}_{i}}(\lambda,v,\{x_{i}\})>0$ for some $i\in\{1,\dots,k\}$, then $\theta_{\mathcal{T}}(\lambda,v,\{\rho\})>0$. In other words, if the contact process dies out on $\mathcal{T}$ then it has to die out on all subtrees $\mathcal{T}_{1},\dots,\mathcal{T}_{k}$. Denote by $g$ the generating function of the random variable $\zeta$. Now we observe that